A=Pe^rt calculate the balance of your account after three years. Round.the answer to the nearest hundredth. Situation you invest 4400 in the account that pays an interest rate of 6.5%, compounded continuously.
Answer by stanbon(75887)

A=Pe^rt calculate the balance of your account after three years. Round.the answer to the nearest hundredth. Situation you invest 4400 in the account that pays an interest rate of 6.5%, compounded continuously.
-------------
A(3) = 4400*e^(0.065*3) = $5347.37
Cheers,
Stan H.
--------------